OpenStudy (anonymous):

2/3(x-6)+2=1/6x+1/2 Need to show work

OpenStudy (anonymous):

3x-6 = 12x + 2 basically you multiply both sides of the equation by 3(x-6). The left side will cancel out and you will have 3x - 6 as the numberator on the right side. Then you do the same for 6x + 1/2. The 6x + 1/2 on the right side will cancel out and you will have (6x + 1/2)*2 on the right side. Now multiply the right side out and you have 3x-6 = 12x + 2. I'll leave that for you to find x.

OpenStudy (anonymous):

junwagh, Good answer but I do not understand how to do this at all..

OpenStudy (anonymous):

http://i.imgur.com/DG74w.png

OpenStudy (dumbcow):

Distribute \[\rightarrow (2/3)x - 4 +2 = (1/6)x +(1/2)\] Combine terms \[(2/3)x - (1/6)x = (1/2) +2\] \[(4/6)x - (1/6)x = (1/2) + (4/2)\] \[(3/6)x = (5/2)\] \[(1/2)x = (5/2)\] Multiply by 2 on both sides \[x = 5\]

OpenStudy (anonymous):

lol i could be reading the problem wrong

OpenStudy (anonymous):

6[2(x-6)/3] + 6(2) = 6[1x/6] + 6(1/2) I multiply everything by six (which is the lowest common denominator. Now solve. 6[(2x-12)/3] + 12 = 6x + 3 2{2x-12) + 12 = 6x + 3 ^I divided the 6 by the 3. 4x-24 +12 = 6x + 3 4x - 6x = 3 +24 - 12 -2x = 13 x = 13/-2 x = -13/2